OpenStudy (18jonea):

Write an equation of a line that passes through (-3, 4) with slope -. @amistre64

OpenStudy (amistre64):

what formula of a line equation implements a slope and a point?

OpenStudy (18jonea):

y-y1= m (x-x1)

OpenStudy (amistre64):

then all we need to do is plug in the values ...

OpenStudy (amistre64):

what do we end up with? and you have not posted a readable m value yet

OpenStudy (18jonea):

-1/2

OpenStudy (18jonea):

is the slope

OpenStudy (amistre64):

ok, then i can verify your results

OpenStudy (18jonea):

y-4= -1/2 (x+3)

OpenStudy (amistre64):

yep, looks good. that is an equation of a line.

OpenStudy (amistre64):

if need be we can rewrite it as y=(stuff) distribute and add thru the 4

OpenStudy (amistre64):

but the question doesnt imply what form is required for a solution.

OpenStudy (18jonea):

y=

OpenStudy (18jonea):

so How would I find the y intercept

OpenStudy (amistre64):

the y axis, is also the x=0 axis ... so the intercept when x=0 is found when we let x=0

OpenStudy (amistre64):

y-4= -1/2 (x+3) y-4= -1/2 (0+3) y-4= -3/2 what is our y value?

OpenStudy (18jonea):

2.5

OpenStudy (18jonea):

so y=-1/2x +2.5

OpenStudy (amistre64):

the other method which i suggested works just the same ... y-4= -1/2 (x+3) y-4= -x/2 -3/2 y= -x/2 -3/2 +4 unless you require decimals, i would keep it all in ratioed numbers y int = 5/2

OpenStudy (amistre64):

not that it matters, its just a visual thing that some teachers adhere to.

OpenStudy (18jonea):

I got it thank you

OpenStudy (amistre64):

:) youre welcome